OpenStudy (anonymous):

The time required to finish a test is normally distributed with a mean of 60 minutes and a standard deviation of 10 minutes. What is the z-Score for a student who finishes the test in 45 minutes?

OpenStudy (anonymous):

I have no clue I would guess 75% but I have no clue

OpenStudy (kropot72):

The z-score is calculated as follows: \[\large z=\frac{X-\mu}{\sigma}=\frac{45-60}{10}=you\ can\ calculate\]